fiND FOUR CONSECUTIVE INTEGERS WHOSE SUM IS 244
Answer by Fombitz(32388)

A+B+C+D=244
Since they are consecutive integers, then
B=A+1
C=B+1=A+2
D=C+1=B+2=A+3
Substituting into the first equation:
A+(A+1)+(A+2)+(A+3)=244
4A+6=244
4A+6-6=244-6 Additive inverse of 6 or (-6)
4A = 238
4A/4 = 238/4
A=238/4
Oops, the results of 238 divided by 4 is not an integer, as you assumed at the beginning. There is no solution.
A quick check: (59,60,61,62) sum to 242 and (60,61,62,63) sum to 246.
No solution.